[Brahms-l] Shift report 20050422 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Apr 22 2005 - 23:58:07 EDT
Shifter leader was JHL
Shifters were 


Beam:
6 Hours beam for physics until 10 pm.

Data:
FS 4deg 1219A 230K Trigger2
FS 4deg 0-Field Run
FS 4deg min-bias run: Trigger4 and Trigger5 only

"Problems":
Some empty H2 channels in monotor. (See e-log)
Some spec. triggered events (~1-2%) have non-event related 
heavy noise in T1. (See e-log)  


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Apr 22 23:58:25 2005

This archive was generated by hypermail 2.1.8 : Fri Apr 22 2005 - 23:58:38 EDT